OpenStudy (anonymous):

the circumference of the bike tire above is 84.78 inches what is the radius of the bike tire [use 3.14 ]

OpenStudy (anonymous):

C = 2*pi*r A= pi*r^2

OpenStudy (anonymous):

Oops, ignore the last equation ^_^